OpenStudy (anonymous):

Find the surface area of the portion of z=4-x^2-y^2 above the xy-plane.

OpenStudy (amistre64):

\[\int\int\sqrt{1+(f_x)^2+(f_y)^2}~dA\]

OpenStudy (amistre64):

z=0 when x^2+y^2 = 4

OpenStudy (amistre64):



OpenStudy (anonymous):

then integrate with respect to x and y?

OpenStudy (amistre64):

f(x,y,z)=4-x^2-y^2 fx=-2x fy=-2y fz = 1, which is already expressed in the general setup

OpenStudy (amistre64):

yes